At what distance would the perspective between two electrons have magnitude 2.00 N? between two protons?


1
Expert's answer
2022-02-05T10:28:20-0500

Protons and electrons have equal charges:


"r=\dfrac e{2\sqrt{\pi\epsilon_0F}}=1.1·10^{-14}\text{ m}."